# Situationally-Aware Dynamics Learning

### v1

@@ V-A1 Simulation:
We use the Gazebo simulator [46]. Specifically, we use Clearpath’s Inspection World [16]. It combines uneven surfaces, slopes, and varying friction properties. These conditions mimic real-world complexities, requiring the robot to adapt dynamically to changes in stability and traction while maintaining precise control, making it an ideal testbed for training and assessing our method.
Furthermore, we create four additional worlds with increasing levels of difficulty. Concretely, these are randomly generated environments created using height maps sampled using Perlin noise [71] that provide a complex and realistic testing ground. These environments feature highly variable and irregular terrain patterns, including sudden changes in elevation, unpredictable slopes, and patches of different roughness and friction. The use of Perlin noise ensures natural-looking, non-repetitive landscapes that challenge the robot’s ability to adapt to dynamic and uneven surfaces. This makes them an effective way to assess the robustness and adaptability of our method in navigating real-world-like conditions. The generated height maps serve as input to the gazebo_terrain package for automatic Gazebo model terrain generation [92]. The parameters used to generate the height maps are presented in Table I.
@@ V-B Baselines
    CaDM: Context-aware Dynamics Model, which aims to learn a “global model that can generalize across different dynamics” [59]. This is the closest baseline to our work, as they also learn a model that can detect and adapt to changes in the underlying dynamics of the agent. Additionally, although they need to learn three models (context encoder, forward and backward dynamics), they do so in parallel instead of relying on a multi-stage curriculum as other recent methods. We use an author-provided and publicly available implementation.
@@
    TD-MPC2: This is one of the leading MBRL methods at the time of writing [37]. It is characterized for learning a latent-space dynamics model, together with an observation encoder, transition reward predictor, terminal value predictor and a return-maximizing action policy. We use an author-provided and publicly available implementation.
@@
    PETS: This method proposed by [15] is a strong state-space dynamics learning method based on a probabilistic ensemble architecture, which in fact inspires our choice of architecture. Our implementation is based on a derivation from MBRL-Lib [72].
@@
    CDL-CMI: This is a causally-inspired, model-based RL method that aims to increase predictive robustness by avoiding unnecessary dependencies between state variables and actions [95]. They also learn a state abstraction for better out-of-distribution predictive performance and improved learning efficiency. We use an author-provided and publicly available implementation.
@@
    Model-Free: We train model-free policies with PPO [78], SAC [34] and TD3 [28]. This is done to present a well-rounded coverage of general RL methods and to get an idea of how recent competitive methods that also rely on auxiliary representations but with model-free policies would compare.
@@
    Dubins: This is a physics-based model [24] that has been used recently as an unstructured terrain navigation baseline in multiple works [85, 102].
@@ V-C3 Software:
All code was developed using ROS Noetic [75], Python 3.10.11 and PyTorch 2.3.0+cu121 [69]. For the sampling-based controllers we use the MBRL-Lib package [72]. In the real-world, for localization we used FasterLIO with IMU preintegration on manifold [6, 101] with the Velodyne 16 LiDAR.
@@ V-D Situationally-Aware Dynamics Learning
Figure 5: Training task in the Inspection World Environment [16]. The region around waypoint #3 is slippery and difficult to reach.
@@
The model $\mathcal{M}_{m}$ is composed of $m=5$ probabilistic neural networks, with three dense layers with 200 neurons each and LeakyReLU non-linearities in between. The output is the mean and log-variance of a Gaussian distribution. Inputs and outputs are normalized using data statistics computed during training. At inference time, predictions are de-normalized with respect to the data statistics obtained during training. We use the Adam optimizer with default hyperparameters [44]. The batch size is 256 with an early-stopping threshold of 1e-3. The model $\mathcal{M}_{m}$ starts with randomly initialized parameters and is trained at the beginning of each episode with the updated replay buffer $\mathcal{R}$, analogous to the MBRL procedure shown in Algorithm 1. A summary of the hyperparameters for the situational awareness module (online hidden state representation learning) and the controller is provided in Table II.
@@ V-E2 Effect of the Hidden State Representations:
We use the Integrated Gradients [88] feature attribution method to verify that our situationally-aware dynamics model relies on the different situations to make its predictions, thus influencing the downstream planner’s control strategies. Figure 10 shows that on average, 60% of the prediction is explained by the current situation the robot is in, while the remaining input features together account for roughly 40%. This clearly highlights that the model is aware of the situation the robot is experiencing and modulates its predictions accordingly.
@@ V-F1 Simulation Experiments:
Control Inconsistency. In Figure 13 we measure the control inconsistency. Formally, it relates to “the error between the robot’s expected navigational behavior and its actual behavior in terms of robot poses” [85]. Our method demonstrates the best performance, followed by the model that incorporates all orientation information (PE-RPY). This result underscores the suitability of our model for sampling-based control. Specifically, our approach strikes the best balance for auto-regressive prediction tasks, such as planning with MPPI. It benefits from the lower one-step prediction error characteristic of more complex models that are harder to learn and generalize, while also offering the reduced prediction space of simpler models (e.g., Dubins), which helps mitigate compounding errors during trajectory rollout.
@@
Table III: Results on four simulation environments with increasing level of difficulty. The robot needs to traverse to fifty randomly generated initial and goal locations in each environment. We evaluate the Success Rate, Position Jerk $[\frac{m}{s^{3}}]$, Orientation Vibration $[\frac{rad}{s}]$ and Orientation Jerk $[\frac{rad}{s^{3}}]$.
@@ A-A Lake Simulator
Defining and identifying the hidden state in the unstructured terrain navigation task is inherently challenging due to the complexity of natural environments and their dynamic properties. Therefore, to further validate our online hidden state representation learning method, we conduct experiments in a controlled environment where we can directly affect the UDGP of the world. For this, we use the Lake World Simulator [9], as shown in Figure 17. This environment provides a practical testbed for analyzing navigation in maritime settings, where dynamic and unpredictable conditions are prevalent. Specifically, we simulate a scenario in which an Autonomous Surface Vehicle (ASV) must navigate across the lake while responding to shifting wave patterns about which it does not have any kind of information beyond what it can measure via its proprioceptive odometry measurements. The ASV’s objective is to navigate toward randomly generated targets while managing to do so with the changing wave patterns.

### v2

@@ V-A1 Simulation:
We use the Gazebo simulator [Koenig-2004-394]. Specifically, we use Clearpath’s Inspection World [inspection_world]. It combines uneven surfaces, slopes, and varying friction properties. These conditions mimic real-world complexities, requiring the robot to adapt dynamically to changes in stability and traction while maintaining precise control, making it an ideal testbed for training and assessing our method.
Furthermore, we create four additional worlds with increasing levels of difficulty. Concretely, these are randomly generated environments created using height maps sampled using Perlin noise [perlin1985image] that provide a complex and realistic testing ground. These environments feature highly variable and irregular terrain patterns, including sudden changes in elevation, unpredictable slopes, and patches of different roughness and friction. The use of Perlin noise ensures natural-looking, non-repetitive landscapes that challenge the robot’s ability to adapt to dynamic and uneven surfaces. This makes them an effective way to assess the robustness and adaptability of our method in navigating real-world-like conditions. The generated height maps serve as input to the gazebo_terrain package for automatic Gazebo model terrain generation [gazebo_terrain]. The parameters used to generate the height maps are presented in Table I.
@@ V-B Baselines
    CaDM: Context-aware Dynamics Model, which aims to learn a “global model that can generalize across different dynamics” [lee2020context]. This is the closest baseline to our work, as they also learn a model that can detect and adapt to changes in the underlying dynamics of the agent. Additionally, although they need to learn three models (context encoder, forward and backward dynamics), they do so in parallel instead of relying on a multi-stage curriculum as other recent methods. We use an author-provided and publicly available implementation.
@@
    TD-MPC2: This is one of the leading MBRL methods at the time of writing [hansen2023td]. It is characterized for learning a latent-space dynamics model, together with an observation encoder, transition reward predictor, terminal value predictor and a return-maximizing action policy. We use an author-provided and publicly available implementation.
@@
    PETS: This method proposed by [chua2018deep] is a strong state-space dynamics learning method based on a probabilistic ensemble architecture, which in fact inspires our choice of architecture. Our implementation is based on a derivation from MBRL-Lib [Pineda2021MBRL].
@@
    CDL-CMI: This is a causally-inspired, model-based RL method that aims to increase predictive robustness by avoiding unnecessary dependencies between state variables and actions [wang2022causal]. They also learn a state abstraction for better out-of-distribution predictive performance and improved learning efficiency. We use an author-provided and publicly available implementation.
@@
    Model-Free: We train model-free policies with PPO [schulman2017proximal], SAC [haarnoja2018soft] and TD3 [fujimoto2018addressing]. This is done to present a well-rounded coverage of general RL methods and to get an idea of how recent competitive methods that also rely on auxiliary representations but with model-free policies would compare.
@@
    Dubins: This is a physics-based model [dubins-dyns] that has been used recently as an unstructured terrain navigation baseline in multiple works [siva2024self, yu2024adaptive].
@@ V-C3 Software:
All code was developed using ROS Noetic [Quigley09], Python 3.10.11 and PyTorch 2.3.0+cu121 [NEURIPS2019_9015]. For the sampling-based controllers we use the MBRL-Lib package [Pineda2021MBRL]. In the real-world, for localization we used FasterLIO with IMU preintegration on manifold [9718203, yu2023fast] with the Velodyne 16 LiDAR.
@@ V-D Situationally-Aware Dynamics Learning
Figure 5: Training task in the Inspection World Environment [inspection_world]. The region around waypoint #3 is slippery and difficult to reach.
@@
The model $\mathcal{M}_{m}$ is composed of $m=5$ probabilistic neural networks, with three dense layers with 200 neurons each and LeakyReLU non-linearities in between. The output is the mean and log-variance of a Gaussian distribution. Inputs and outputs are normalized using data statistics computed during training. At inference time, predictions are de-normalized with respect to the data statistics obtained during training. We use the Adam optimizer with default hyperparameters [kingma2014adam]. The batch size is 256 with an early-stopping threshold of 1e-3. The model $\mathcal{M}_{m}$ starts with randomly initialized parameters and is trained at the beginning of each episode with the updated replay buffer $\mathcal{R}$, analogous to the MBRL procedure shown in Algorithm 1. A summary of the hyperparameters for the situational awareness module (online hidden state representation learning) and the controller is provided in Table II.
@@ V-E2 Effect of the Hidden State Representations:
We use the Integrated Gradients [integrated_gradients] feature attribution method to verify that our situationally-aware dynamics model relies on the different situations to make its predictions, thus influencing the downstream planner’s control strategies. Figure 10 shows that on average, 60% of the prediction is explained by the current situation the robot is in, while the remaining input features together account for roughly 40%. This clearly highlights that the model is aware of the situation the robot is experiencing and modulates its predictions accordingly.
@@ V-F1 Simulation Experiments:
Control Inconsistency. In Figure 13 we measure the control inconsistency. Formally, it relates to “the error between the robot’s expected navigational behavior and its actual behavior in terms of robot poses” [siva2024self]. Our method demonstrates the best performance, followed by the model that incorporates all orientation information (PE-RPY). This result underscores the suitability of our model for sampling-based control. Specifically, our approach strikes the best balance for auto-regressive prediction tasks, such as planning with MPPI. It benefits from the lower one-step prediction error characteristic of more complex models that are harder to learn and generalize, while also offering the reduced prediction space of simpler models (e.g., Dubins), which helps mitigate compounding errors during trajectory rollout.
@@
Table III: Results on four simulation environments with increasing level of difficulty. The robot needs to traverse to fifty randomly generated initial and goal locations in each environment. We evaluate the Success Rate, Position Jerk $\frac{m}{s^{3}}]$, Orientation Vibration $\frac{rad}{s}]$ and Orientation Jerk $\frac{rad}{s^{3}}]$.
@@ A-A Lake Simulator
Defining and identifying the hidden state in the unstructured terrain navigation task is inherently challenging due to the complexity of natural environments and their dynamic properties. Therefore, to further validate our online hidden state representation learning method, we conduct experiments in a controlled environment where we can directly affect the UDGP of the world. For this, we use the Lake World Simulator [bingham19toward], as shown in Figure 17. This environment provides a practical testbed for analyzing navigation in maritime settings, where dynamic and unpredictable conditions are prevalent. Specifically, we simulate a scenario in which an Autonomous Surface Vehicle (ASV) must navigate across the lake while responding to shifting wave patterns about which it does not have any kind of information beyond what it can measure via its proprioceptive odometry measurements. The ASV’s objective is to navigate toward randomly generated targets while managing to do so with the changing wave patterns.